# Multidata Causal Discovery for Statistical Hurricane Intensity Forecasting

### v1

@@ 4.2 Causal feature selection outperforms baseline methods
Figure 3: Summary of results for the 24-hour intensity change forecast (DELV24) using SHIPS+ERA5 predictors without SHIPS link assumptions. (a) Bar plot showing the frequency of each variable’s selection across the best models from all seven cross-validation folds. A red dashed line marks the threshold (more than 3 folds) used to shortlist robust predictors for inclusion in the final SHIPS+ list. (b) Boxplot comparing test R² values for target DELV for each lead times 24, 48, 72, 96, 120 hrs for experiments with kitchen-sink approach (without link assumptions) across four feature selection strategies: causal discovery, correlation ranking, random forest importance, and no selection. Causal feature selection yields the highest median R² until 72 hrs lead time, showing improved generalization in a purely statistical prediction setup.
@@
Figure 3 summarizes the process and outcomes of identifying robust predictors using our causal feature selection pipeline for the 24-hour intensity change forecast (DELV24). Panel (a) shows how often each candidate predictor was selected in the seven cross-validation folds. By applying a clear threshold, which requires a predictor to appear in more than three folds—we filter out variables with inconsistent contributions and focus on predictors with stable links to intensity change. The color gradient highlights the strength of each variable’s presence, from vivid cyan for consistently selected predictors to vivid pink for those with minimal presence. Panel (b) compares test $R^{2}$ scores for DELV24 across four feature selection strategies. Causal feature selection achieves the highest median skill, outperforming correlation ranking, random forest importance, and the no-selection baseline. This demonstrates that incorporating causally motivated predictors yields clear improvements in model generalizability. Comparing the causal and correlation feature selection methods, we observe that causal yields fewer outliers in prediction errors compared to the correlation-based approach, yet displays slightly larger uncertainty bounds. This behavior can be attributed to the fact that PC-based causal discovery method tends to identify stable, interventionally relevant relationships rather than purely predictive ones. When applied to a system with underlying nonlinear dynamics, the PC-based model may underfit complex regions of the predictor space, leading to broader prediction intervals, which reflects the model’s caution in extrapolating beyond what the causal structure supports. This caution mitigates against large, spurious prediction errors, and results in fewer outliers and more robust predictions on unseen test cases. The trade-off between model stability and uncertainty shown here highlights the value of causal methods over correlation for building reliable statistical models for high-impact forecasting applications. By extending this comparison across all lead times (Fig S11, S12 in the SI) and evaluating the consistency of variable selection, we ultimately created a shortlist a new set of predictors to be added to SHIPS. These variables reliably contribute to improved forecast skill for short-range intensity prediction, particularly up to 72 hours. Beyond this time frame, the benefit of causal feature selection diminishes, emphasizing that while our method provides meaningful gains for in short-range, longer lead times will likely require integrating statistical models with dynamical forecast guidance to capture additional sources of predictability.
@@ 4.3 Recommending causally relevant predictors
Here, we use the coefficient of determination ($R^{2}$) as the primary performance metric, where positive values indicate that the model explains some fraction of the variance in TC intensity change, while negative values indicate that the model performs worse than simply predicting the mean. To increase the robustness of the predictor list and reduce the uncertainties arising from model hyperparameters and cross-validation strategies sweet2023cross , the variable lists from the best MLR models for different folds are aggregated as a summary variable list. Of the different variables in the list, only those that are chosen in more than half of the cross-validation folds (more than 3 times) will be considered as candidate features to be included in the operational SHIPS for testing. Comparison of frequently repeated variables for experiments with and without SHIPS link assumptions (Fig. 3 a, Fig. S11, Fig. S12) yields a final shortlist of six final predictors that mostly describe lower and middle tropospheric vertical wind shear (e.g. SHL0, SHL1, SHMD), surface and boundary layer moisture (R000, R001), and midtropospheric potential vorticity conditions (PVOR) in the outer area of TC. The complete list of additional predictors recommended is provided in Table 1.
@@
Before putting the discovered variables into operational models for testing, it is important to ensure that they pass the “expert judgment”, as there is reasonable plausibility that these predictors could be causally related to TC intensity change. Low-tropospheric wind shear has been shown to be more negatively correlated with the intensity of Pacific typhoons that occurred in active typhoon seasons than commonly used deep-layer shear wang2015statistical . fu2019effect further conclude that low-level shear produces quasiperiodic oscillations in the intensity of the TC that are related to the variability in the moisture of the boundary layer induced by the TC rainbands. Although TC intensification is generally most associated with relative humidity above the boundary layer wu2012relationship , the moisture above the boundary layer reflects the balance between surface flux moistening and dry air ventilation near the surface fischer2023tale . The 1000 hPa relative humidity in the outer environment (R000), is a variable available in the SHIPS developmental file but unused in operational models. This shows that causal selection can reveal relevant predictors missed by traditional screening, improving short-range skill. This suggests that addition of overlooked factors like surface RH alongside mid-tropospheric humidity and shear helps capture multiscale processes drive intensity change, supporting the idea that boundary layer moisture could play a key role in regulating TC intensities wadler2021thermodynamic and the onset of rapid intensification chen2021boundary . Finally, mid-tropospheric vortex strength is critical in the early intensification of TCs bister1997genesis ; yang2020interactive .
@@ 4.4 Added value of the nonlinear SHIPS+ models
Figure 5: Predictor importance and dependencies for models trained on SHIPS+. (a–b) Global feature importance ranked by mean $\left|\textrm{SHAP}\right|$ for the MLP (a) and MLR (b). (c) SHAP dependence for baseline SHIPS predictors POT (potential intensity minus current intensity) and T200 (200-hPa temperature, 200–800 km-averaged). (d) SHAP dependence for causally selected predictors SHL1 (1000–850-hPa vertical wind shear, 200–1000 km-averaged) and PVOR (500-hPa potential vorticity, 200–800 km-averaged). (e) For near-surface humidity, the MLP learns opposite dependencies: negative with R001 (1000-hPa relative humidity, 0–500 km-averaged) and positive with R000 (1000-hPa relative humidity, 200–800 km-averaged). All predictors in panels c-e are standardized.
@@
To better understand why the nonlinear SHIPS+ models outperform the other models, we conducted a SHAP (SHapley Additive exPlanations) analysis for the 24-hr lead time MLP and MLR models using both SHIPS and SHIPS+ datasets (see Fig. 5). SHAP values provide a model-agnostic interpretation of feature importance by quantifying the marginal contribution of each predictor to the model’s output. To conduct the SHAP analysis, we rely on Kernel SHAP kernel_shap_2017 , using 300 samples from the training dataset to retrieve the background signal and evaluating on the 240 samples in the test set. Focusing on the differences between the linear MLR and the MLP models, we observe that of the six causal predictors, the MLP primarily utilizes the near-surface moisture predictors (red bars in Fig.5a), whereas the SHIPS+ MLRs (Fig.5b) use the causal shear predictors instead. A potential reason for MLPs and MLRs to select different causal predictors is that the relationship between near-surface moisture and short-term changes in TC intensity is too nonlinear to be used by linear MLRs. The dependencies of the causal boundary layer moisture predictors (e.g., R001) on the MLR predictions are nonlinear (Fig. 5e). Apart from prioritizing additional nonlinear dependencies, the SHIPS+ MLPs could also overperform their MLR counterparts by learning different dependencies for the existing SHIPS predictors. Figure 5d presents two examples in which this is the case, where low-level wind shear (SHL1) and mid-tropospheric potential vorticity (PVOR) have a much stronger dependence on MLP predictions than the muted ones for MLR models. Furthermore, MLPs learned a dependence between SHL1 and short-term intensity evolution that is strikingly different from the MLR dependence. Although intensification predictions in both the nonlinear and linear regression frameworks have a similar dependency on potential intensity (POT), which is the most critical predictor in both types of models, the difference in the selection of other leading predictors and their respective learned dependencies contribute to the significant differences in the generalizability of the model to test TCs (Fig.  4).
@@ 4.5 Linear vs Nonlinear Models with Causal Predictors: Case Study
To illustrate the role of causally relevant predictors, we present Hurricane Larry (2021) as a case study.We selected Larry because, among the storms in the independent test set, it showed the largest increase in predictive skill (R²) when using SHIPS+ MLPs with causal predictors compared to the operational SHIPS model. This allows us to investigate which features and causal predictors contributed most to the improvement in forecast performance. Larry was a long-lived and intense Cape Verde hurricane that underwent a period of rapid intensification over the tropical Atlantic before eventually making landfall in Newfoundland as a Category 1 hurricane. For this analysis, we truncate the time series before its extratropical transition, focusing solely on the tropical phase. Figure  6a) shows the best track of Larry during the evaluation period (Left), and a comparison of model performance (Right) between MLR and MLP, both with and without the addition of causally selected predictors. MLR SHIPS and MLR SHIPS+ both generally under-predict intensity change throughout the period, showing less variability and weaker response to fluctuations. Both versions of the Multi-Layer Perceptron (MLP) models with SHIPS and SHIPS+ (causally selected features) are consistently closer to the ground truth (black line) than the corresponding MLR models.
Figure 6: Hurricane Larry (2021): (a) Best track from IBTrACS showing the time used for testing (b) performance comparison of MLR vs. MLP comparing IBTrACS 24 hr intensity change (truth) to predictions with (SHIPS+) and without causal predictors (SHIPS). Overall, the MLP consistently outperforms the MLR, with the addition of predictors improving the performance notable from 54 hours. (c) Lifetime-maximum absolute SHAP values highlighting the dominant role of the added predictors (R001, PVOR). (d) Time-series decomposition of $\Delta y_{\mathrm{pred}}$ showing R001 and PVOR as the main contributors, with smaller adjustments from LAT and POT.
MLP SHIPS initially over-predicts, but adding SHIPS+ dampens this overestimation, bringing it closer to the true signal, especially after 54 hours. In contrast, the performance gain from adding causal variables to MLR is minimal, indicating that linear models cannot fully leverage the additional causal features, highlighting the limitations of linear regression for dynamic systems like TCs, while the MLP models show a more dynamic response to intensity fluctuations.
To assess the role of causal predictors, we test whether the SHIPS++ MLP’s gains arise from learning new relations for the added predictors or simply from reweighting existing SHIPS predictors. Therefore, we apply the SHAP-based decomposition of grundner2021deep (Appendix B) to compare the SHIPS model $f$ and the SHIPS++ model $g$. By SHAP additivity, the prediction difference can be decomposed as follows:
@@
For Hurricane Larry, lifetime-mean absolute SHAP values (Fig.6c) indicate that Eq.3 is well approximated by the contributions of two added predictors (R001 and PVOR) plus adjustments to two existing predictors (LAT and POT). The time-series decomposition (Fig.6d) confirms that R001 and PVOR dominate $\Delta y_{\mathrm{pred}}$, with smaller adjustments from LAT and POT and a small residual from all other features over the period of interest (30-120 hr). This lends credence to the interpretation that SHIPS++ improvements primarily arise from the added causal predictors. Consistent with the partial dependence for R001 (Fig.5e), SHIPS overestimates Larry’s intensification because the original predictors under-represent the suppressing influence of inner-domain near-surface humidity; including R001 mitigates this bias.
@@ 4.6 Testing the new predictors in operational SHIPS
As described by demaria2022national , many new predictors have been added to SHIPS since the operational version was first implemented in 1991. As a preliminary test of how the research results presented above might improve the operational SHIPS model, the six potential new predictors listed in Table 1 were evaluated using the standard SHIPS procedures for annual updates. For this test, the North Atlantic data from 1982-2021 were used for training and the 2022-2024 cases were used for testing. This procedure does not use validation data because the SHIPS prediction coefficients are uniquely determined from the MLR fit to the training data. SHIPS uses GFS model fields so the predictors in Table 1 were recalculated from the GFS and added to the 28 predictors in the 2025 operational SHIPS model.
@@
As in SHIPS development, candidate variables are added to the developmental set and undergo a predictor-screening test for statistical significance with respect to intensity changes at 6-h increments to lead times of 168 h. Training follows the SHIPS perfect-prognosis protocol: predictors are averaged over each forecast interval and sampled from GFS analysis fields about the future best-track positions at the valid times. In operations, the same predictors are computed from GFS forecast fields along the National Hurricane Center (NHC) forecast track. Note that the causal predictor identification used in this manuscript (Table 1) differs by using only $t=0$ values about the current best-track position—by design, to keep feature discovery in a purely statistical forecasting (not post-processing) setting and to make the causal tests consistent with the regression task. We then apply the standard SHIPS screening and coefficient-update procedure described next. Three conditions are needed to pass the screening step: (1) When added one at a time, a new predictor must increase the variance explained by the model by at least 0.2% averaged over 5 consecutive forecast intervals; (2) The regression coefficient must be significant at the 99% level for at least 5 forecast intervals; (3) All predictors that pass steps 1 and 2 are added to the training sample and the coefficients are recalculated. The regression coefficients for each predictor must still be significant at the 99% level for at least five forecast intervals. If a predictor does not pass step 3, the least significant predictor is removed and then step 3 is repeated until all retained predictors pass the significance test. Results showed that SHMD, R001 and PVOR passed steps 1 and 2 for the 1982-2021 North Atlantic training sample. When all three were added, they also passed step 3. The next step in testing new SHIPS predictors is to perform retrospective forecasts on independent cases using only the input that is available in real time (forecast tracks and GFS forecast fields). To allow for a fair comparison, SHIPS was trained on the 1982-2021 sample with the original 28 operational predictors (baseline) and then with the addition of the three new predictors and both versions were run on the 2022-2024 independent cases. This sample included 794 cases with a 12 h forecast, which decreased to 114 cases by 168 h since many TCs dissipate by seven days. The mean absolute error (MAE) of the intensity forecasts for the baseline SHIPS and with the three new predictors were calculated using the intensity in the final NHC best track as “truth” following the standard NHC forecast procedure CangialosiReinhartMartinez2024 . The verification sample includes tropical and subtropical cyclones but excludes the extratropical and pre-genesis stages.
@@
Figure 7 shows the percent improvement (reduction in MAE) of SHIPS with the new predictors relative to baseline. SHIPS forecasts improved at all forecast times, with the largest improvements at longer forecast times. The improvements at 120-168 h were statistically significant at the 90% level using a standard statistical test that accounts for serial correlation. Although causal discovery used only $t=0$ values, the longer-lead time gains seen in Figure 7 likely reflect the statistical–dynamical use of forecast fields in operations, motivating extension of the discovery procedure to predictors evaluated during the forecast period. The three new predictors (SHMD, R001 and PVOR) will be considered for implementation in future operational versions of SHIPS.

### v2

@@ 4.2 Causal feature selection outperforms baseline methods
Figure 4: Summary of results for the 24 hour intensity change forecast (DELV24) using SHIPS+ERA5 predictors for the noASSUMPS experiment. (a) Bar plot showing the frequency of each variable’s selection across the best models from all seven cross-validation folds. A red dashed line marks the threshold (more than 3 folds) used to shortlist robust predictors for inclusion in the final SHIPS+ list. (b) Boxplot comparing test R² values for target DELV for each lead times 24, 48, 72, 96, 120 hrs for experiments with kitchen-sink approach (without link assumptions) across four feature selection strategies: causal discovery, correlation ranking, random forest importance, and no selection. Causal feature selection yields the highest median R² until 72 hrs lead time, showing improved generalization in a purely statistical prediction setup.
@@
Figure 4 summarizes the process and outcomes of identifying robust predictors using our causal feature selection pipeline for the 24 hour intensity change forecast (DELV24). Panel (a) shows how often each candidate predictor was selected in the seven cross-validation folds. By applying a clear threshold, which requires a predictor to appear in more than three folds—we filter out variables with inconsistent contributions and focus on predictors with stable links to intensity change. Panel (b) compares test $R^{2}$ scores for DELV24 across four feature selection strategies. Causal feature selection achieves the highest median skill, outperforming correlation ranking, random forest importance, and the no-selection baseline. This demonstrates that incorporating causally motivated predictors yields clear improvements in model generalizability. Comparing the causal and correlation feature selection methods, we observe that causal yields fewer outliers in prediction errors compared to the correlation-based approach, yet displays slightly larger uncertainty bounds. This behavior can be attributed to the fact that PC-based causal discovery method tends to identify stable, interventionally relevant relationships rather than purely predictive ones. When applied to a system with underlying nonlinear dynamics, the PC-based model may underfit complex regions of the predictor space, leading to broader prediction intervals, which reflects the model’s caution in extrapolating beyond what the causal structure supports. This caution mitigates against large, spurious prediction errors, and results in fewer outliers and more robust predictions on unseen test cases. The trade-off between model stability and uncertainty shown here highlights the value of causal methods over correlation for building reliable statistical models for high-impact forecasting applications. By extending this comparison across all lead times (Fig S11, S12 in the SI) and evaluating the consistency of variable selection, we ultimately shortlisted six predictors to be added to SHIPS. These variables reliably contribute to improved forecast skill for short-range intensity change prediction, particularly up to 72 hours. Beyond this time frame, the benefit of causal feature selection diminishes, emphasizing that while our method provides meaningful gains for in short-range, longer lead times will likely require integrating statistical models with dynamical forecast guidance to capture additional sources of predictability.
@@ 4.3 Recommending causally relevant predictors
Here, we use the coefficient of determination ($R^{2}$) as the primary performance metric, where positive values indicate that the model explains some fraction of the variance in TC intensity change, while negative values indicate that the model performs worse than simply predicting the mean. To increase the robustness of the predictor list and reduce the uncertainties arising from model hyperparameters and cross-validation strategies Sweet et al. (2023), the variable lists from the best MLR models for different folds are aggregated as a summary variable list. Of the different variables in the list, only those that are chosen in more than half of the cross-validation folds (at least 4 times) will be considered as candidate features to be included in the operational SHIPS for testing. Applying this shortlisting criterion and comparing frequently repeated variables across experiments withASSUMPS and noASSUMPS (Fig. 4 a, Fig. S11, Fig. S12) yields a final shortlist of six final predictors that mostly describe lower and middle tropospheric vertical wind shear (e.g. SHL0, SHL1, SHMD), surface and boundary layer moisture (R000, R001), and midtropospheric potential vorticity conditions (PVOR) in the outer area of TC. The complete list of additional predictors recommended is provided in Table 1. The most frequently selected causal predictors describe the kinematic and thermodynamic environment in the TC outer core, which likely reflects the critical implication of TC outer rainbands on TC structure and intensity Wang (2009); Yu et al. (2021).
@@
Before testing discovered variables in operational models, it is important to provide evidence that they are physically related to TC intensity change based on current knowledge. Low-tropospheric wind shear, which usually promotes linear convective organization rather than circular organization Robe and Emanuel (2001); Finocchio et al. (2016), is more negatively correlated with the intensity of Pacific typhoons that occurred in active typhoon seasons than commonly used deep-layer shear Wang et al. (2015).Low-level shear is further concluded to produce quasiperiodic oscillations in TC intensity that are related to variability in boundary-layer moisture induced by TC rainbands Fu et al. (2019). TC intensification is generally associated with relative humidity above the boundary layer Wu et al. (2012), as reflected by the inclusion of RHMD in the original SHIPS predictors. In contrast, the 1000 hPa relative humidity in the outer environment (R000) is a variable unused in operational models but available in the extended predictor list and selected by causal discovery. This selection can be physically explained as moisture within the boundary layer reflects the balance between surface flux moistening and dry air ventilation at the top of the boundary layer, through downdraft ventilation in sheared, dry environments Alland et al. (2021), which can strongly modulate convective buoyancy and intensification rates Riemer et al. (2010); Fischer et al. (2023). This shows that causal selection can reveal relevant predictors missed by traditional screening, improving short-range skills. This suggests that addition of overlooked factors like surface relative humidity alongside mid-tropospheric humidity and shear helps capture multiscale processes drive intensity change, supporting the idea that boundary layer moisture could play a key role in regulating TC intensities Wadler et al. (2021) and the onset of RI Chen et al. (2021). Finally, mid-tropospheric vortex strength is critical in the early intensification of TCs Bister and Emanuel (1997); Yang and Tan (2020).
@@ 4.4 Added value of the nonlinear SHIPS+ models
Figure 6: Predictor importance and dependencies for models trained on SHIPS+. (a–b) Global feature importance ranked by mean $\left|\textrm{SHAP}\right|$ for the MLP (a) and MLR (b). (c) SHAP dependence for baseline SHIPS predictors POT (potential intensity minus current intensity) and T200 (200 hPa temperature, 200–800 km-averaged). (d) SHAP dependence for causally selected predictors SHL1 (1000–850 hPa vertical wind shear, 200–1000 km-averaged) and PVOR (500 hPa potential vorticity, 200–800 km-averaged). (e) For near-surface humidity, the MLP learns opposite dependencies: negative with R001 (1000 hPa relative humidity, 0–500 km-averaged) and positive with R000 (1000 hPa relative humidity, 200–800 km-averaged). All predictors in panels c-e are standardized.
@@
To better understand why the nonlinear SHIPS+ models outperform the other models, we conducted a SHAP (SHapley Additive exPlanations) analysis for the 24 hour lead time MLP and MLR models using both SHIPS and SHIPS+ datasets (see Fig. 6). SHAP values provide a model-agnostic interpretation of feature importance by quantifying the marginal contribution of each predictor to the model’s output. To conduct the SHAP analysis, we rely on Kernel SHAP Lundberg and Lee (2017), using 300 samples from the training dataset to retrieve the background signal and evaluating on the 240 samples in the test set. Focusing on the differences between the linear MLR and the MLP models, we observe that of the six causal predictors, the MLP primarily utilizes the near-surface moisture predictors (red bars in Fig.6a), whereas the SHIPS+ MLRs (Fig.6b) use the causal shear predictors instead. A potential reason for MLPs and MLRs to select different causal predictors is that the relationship between near-surface moisture and short-term changes in TC intensity is too nonlinear to be used by linear MLRs. The dependencies of the causal boundary layer moisture predictors (e.g., R001) on the MLR predictions are nonlinear (Fig. 6e). Interestingly, the SHAP dependence plots reveal that the MLP learns opposing relationships for inner-core and outer-core near-surface humidity (Fig. 6e): increases in R001 (0–500 km) are associated with weaker predicted intensification, whereas increases in R000 (200–800 km) support stronger intensification. This contrast indicates that the MLP is leveraging radial gradients in boundary-layer moisture, a signal that the linear MLR framework does not capture. Apart from prioritizing additional nonlinear dependencies, the SHIPS+ MLPs could also overperform their MLR counterparts by learning different dependencies for the existing SHIPS predictors. Figure 6d presents two examples in which this is the case, where low-level wind shear (SHL1) and mid-tropospheric potential vorticity (PVOR) have a much stronger dependence on MLP predictions than the muted ones for MLR models. The negative correlation between intensification and PVOR might reflect the negative impact of TC outer rainbands on intensification Wang (2009); Yu et al. (2021). Furthermore, MLPs learned a dependence between SHL1 and short-term intensity evolution that is strikingly different from the MLR dependence. Although intensification predictions in both the nonlinear and linear regression frameworks have a similar dependency on potential intensity (POT), which is the most critical predictor in both types of models, the difference in the selection of other leading predictors and their respective learned dependencies contribute to the significant differences in the generalizability of the model to test TCs (Fig.  5).
@@ 4.5 Linear vs Nonlinear Models with Causal Predictors: Case Study
To illustrate the role of causally relevant predictors, we present Hurricane Larry (2021) as a case study. We selected Larry because, among the storms in the independent test set, it showed the largest increase in predictive skill (R²) when using SHIPS+ MLPs with causal predictors compared to the operational SHIPS model. This allows us to investigate which features and causal predictors contributed most to the improvement in forecast performance. Larry was a long-lived and intense Cape Verde hurricane that underwent a period of RI over the tropical Atlantic before eventually making landfall in Newfoundland as a Category 1 hurricane. For this analysis, we truncate the time series before its extratropical transition, focusing solely on the tropical phase. Figure 7a shows the best track of Larry during the evaluation period (Left), and a comparison of model performance (Right) between MLR and MLP, both with and without the addition of causally selected predictors. MLR SHIPS and MLR SHIPS+ both generally under-predict intensity change throughout the period, showing less variability and weaker response to fluctuations. Both versions of the Multi-Layer Perceptron (MLP) models with SHIPS and SHIPS+ (causally selected features) are consistently closer to the ground truth (black line) than the corresponding MLR models.
Figure 7: Hurricane Larry (2021): (a) Best track from IBTrACS showing the time used for testing (b) performance comparison of MLR vs. MLP comparing IBTrACS 24 hour intensity change (truth) to predictions with (SHIPS+) and without causal predictors (SHIPS). Overall, the MLP consistently outperforms the MLR, with the addition of predictors improving the performance notable from 54 hours. (c) Lifetime-maximum absolute SHAP values highlighting the dominant role of the added predictors (R001, PVOR). (d) Time-series decomposition of $\Delta y_{\mathrm{pred}}$ showing R001 and PVOR as the main contributors, with smaller adjustments from LAT and POT.
MLP SHIPS generally over-predicts TC Larry’s intensification, except between Hour 24-48. Adding SHIPS+ dampens this overestimation, bringing it closer to the true signal, especially after 54 hours. In contrast, the performance gain from adding causal variables to MLR is minimal, indicating that linear models cannot fully leverage the additional causal features, highlighting the limitations of linear regression for dynamic systems like TCs, while the MLP models show a more dynamic response to intensity fluctuations.
To assess the role of causal predictors, we test whether the SHIPS++ MLP’s gains arise from learning new relations for the added predictors or simply from reweighting existing SHIPS predictors. Therefore, we apply the SHAP-based decomposition of Grundner et al. (2022) (Appendix B) to compare the SHIPS model $f$ and the SHIPS++ model $g$. By SHAP additivity, the prediction difference can be decomposed as follows:
@@
For Hurricane Larry, lifetime-mean absolute SHAP values (Fig.7c) indicate that Eq.3 is well approximated by the contributions of two added predictors (R001 and PVOR) plus adjustments to two existing predictors (LAT and POT). The time-series decomposition (Fig.7d) confirms that R001 and PVOR are the two most important contributors to $\Delta y_{\mathrm{pred}}$, with smaller adjustments from LAT and POT and a small residual from all other features over the period of interest (30-120 hour). This lends credence to the interpretation that SHIPS++ improvements primarily arise from the added causal predictors. Consistent with the partial dependence for R001 (Fig.6e), SHIPS underestimates the weakening of Larry after Hour 54 because the original predictors under-represent the suppressing influence of inner-domain near-surface humidity; including R001 mitigates this bias.
@@ 4.6 Testing the new predictors in operational SHIPS
As described by DeMaria et al. (2022), many new predictors have been added to SHIPS since the operational version was first implemented in 1991. As a preliminary test of how the research results presented above might improve the operational SHIPS model, the six potential new predictors listed in Table 1 were evaluated using the standard SHIPS procedures for annual updates. For this test, the North Atlantic data from 1982-2021 were used for training and the 2022-2024 cases were used for testing. This procedure does not use validation data because the SHIPS prediction coefficients are uniquely determined from the MLR fit to the training data. SHIPS uses GFS model fields so the predictors in Table 1 were recalculated from the GFS and added to the 28 predictors in the 2025 operational SHIPS model.
@@
As in SHIPS development, candidate variables are added to the developmental set and undergo a predictor-screening test for statistical significance with respect to intensity changes at 6 hour increments to lead times of 168 h. Training follows the SHIPS perfect-prognosis protocol: predictors are averaged over each forecast interval and sampled from GFS analysis fields about the future best-track positions at the valid times. In operations, the same predictors are computed from GFS forecast fields along the National Hurricane Center (NHC) forecast track. Note that the causal predictor identification used in this manuscript (Table 1) differs by using only $t=0$ values about the current best-track position—by design, to keep feature discovery in a purely statistical forecasting (not post-processing) setting and to make the causal tests consistent with the regression task. We then apply the standard SHIPS screening and coefficient-update procedure described next. Three conditions are needed to pass the screening step: (1) When added one at a time, a new predictor must increase the variance explained by the model by at least 0.2% averaged over 5 consecutive forecast intervals; (2) The regression coefficient must be significant at the 99% level for at least 5 forecast intervals; (3) All predictors that pass steps 1 and 2 are added to the training sample and the coefficients are recalculated. The regression coefficients for each predictor must still be significant at the 99% level for at least five forecast intervals. If a predictor does not pass step 3, the least significant predictor is removed and then step 3 is repeated until all retained predictors pass the significance test. Results showed that SHMD, R001 and PVOR passed steps 1 and 2 for the 1982-2021 North Atlantic training sample. When all three were added, they also passed step 3. The next step in testing new SHIPS predictors is to perform retrospective forecasts on independent cases using only the input that is available in real time (forecast tracks and GFS forecast fields). To allow for a fair comparison, SHIPS was trained on the 1982-2021 sample with the original 28 operational predictors (baseline) and then with the addition of the three new predictors and both versions were run on the 2022-2024 independent cases. This sample included 794 cases with a 12 hour forecast, which decreased to 114 cases by 168 hours since many TCs dissipate by seven days. The mean absolute error (MAE) of the intensity forecasts for the baseline SHIPS and with the three new predictors were calculated using the intensity in the final NHC best track as “truth” following the standard NHC forecast procedure Cangialosi et al. (2024). The verification sample includes tropical and subtropical cyclones but excludes the extratropical and pre-genesis stages.
@@
Figure 8 shows the percent improvement (reduction in MAE) of SHIPS with the new predictors relative to baseline. SHIPS forecasts improved at all forecast times, with the largest improvements at longer forecast times. The relatively small improvements at short lead times reflect the dominant role of the persistence term in operational SHIPS, together with strong constraints imposed by initial conditions and directly assimilated observations, which limit the contribution of GFS-based environmental predictors. As forecast lead time increases, the influence of persistence and initial conditions weakens, while large-scale environmental controls become more important, allowing the newly identified predictors to produce larger gains in forecast skill. The improvements at 120-168 hours were statistically significant at the 90% level using a standard statistical test that accounts for serial correlation. Although causal discovery used only $t=0$ values, the longer-lead time gains seen in Figure 8 likely reflect the statistical–dynamical use of forecast fields in operations, motivating extension of the discovery procedure to predictors evaluated during the forecast period. The three new predictors (SHMD, R001 and PVOR) will be considered for implementation in future operational versions of SHIPS.